# Understanding Optimal Feature Transfer via a Fine-Grained Bias-Variance Analysis

### v1

@@ 3 Analytic Results for Downstream Risk
3.1 Fine-Grained Bias-Variance Decomposition
The semi-classical (SC) bias-variance decomposition of the risk $\mathbb{E}_{{\bm{\varepsilon}}}R$ is defined as $B_{\mathsf{SC}}+V_{\mathsf{SC}}$ with
@@ 3.1 Fine-Grained Bias-Variance Decomposition
for new response ${y}_{\mathrm{new}}={\mathbf{x}_{\mathrm{new}}}^{\top}{{\bm{\beta}^{\star}}}$. The bias-variance decomposition above is widely recognized in the statistics literature [27].
However, as pointed out in [4], the above is conditional on $\mathbf{X}$ and only decomposes randomness in label noise ${\bm{\varepsilon}}$. Following [4], we focus on the fine-grained (FG) bias-variance decomposition
@@
where $B:=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\left({y}_{\mathrm{new}}-\mathbb{E}_{\mathbf{X}\!\!,{\bm{\varepsilon}}}\hat{y}\right)^{2}$ and $V:=V_{\mathbf{X}}+V_{\mathbf{X},{\bm{\varepsilon}}}+V_{{\bm{\varepsilon}}}$ with
@@
Here, $B$ and $V$ are the bias and variance with respect to randomness in both the data and label noise; $V_{\mathbf{X}}$, $V_{{\bm{\varepsilon}}}$, and $V_{\mathbf{X},{\bm{\varepsilon}}}$ are a two-way ANOVA decomposition of $V$: $V_{\mathbf{X}}$ and $V_{{\bm{\varepsilon}}}$ are the variances explained by data and noise individually, and $V_{\mathbf{X},{\bm{\varepsilon}}}$ the additional variance explained by data and noise jointly. Note the relationships
between the SC and FG decompositions.
We now explain the benefits of considering the fine-grained bias-variance decomposition for this paper: (i) unlike the SC decomposition, FG decomposition provides a clear interpretation of each additive component within the asymptotic risk formula (see Equation 14 and discussion below); (ii) interpreting the SC bias can be challenging, especially as it tends to diverge at the interpolation boundary. The FG decomposition resolves this by attributing the divergence to the variance component $V_{\mathbf{X}}$, while the FG bias $B$ exhibits straightforward monotonic trends in our ablation studies (see Figure 1 and 2); (iii) theoretical and empirical analysis of the FG-decomposed components uncovers a new phase transition within the $n>p$ regime (see Section 5).
@@ 3.2 Exact Asymptotics
Define $H:=\{i\in\{1,\ldots,p\}:\hat{t}_{i}\neq 0\}$, where |$\absolutevalue{H}=h$. Let $b_{0}\in\mathbb{R}_{+}$ be the unique non-negative solution of
@@
The following theorem states that in the sample-deficient regime ($n<h=\mathrm{rank}(\mathbf{\Sigma})$), the total risk $R$ and its FG-decomposed components $B,V_{\mathbf{X}},V_{\mathbf{X},{\bm{\varepsilon}}}$ converge to $\mathfrak{R}$ and $\mathfrak{B},\mathcal{V}\mathfrak{B}$ and $\mathcal{V}$ respectively whereas in the sample-rich regime ($n>h=\mathrm{rank}(\mathbf{\Sigma})$), the total risk has only one non-zero component $V_{\mathbf{X},{\bm{\varepsilon}}}$ and it converges to $\mathcal{U}$. We defer the proof to Appendix B.
@@
Remark 3.4 (Technical Novelty).
Note that by (8), our setting reduces to the standard setting of plain ridgeless regression with feature covariance $\hat{\bm{\Gamma}}^{-1/2}\mathbf{\Sigma}\hat{\bm{\Gamma}}^{-1/2}$. The latter setting has been explored in numerous prior works [18, 26, 14] using random matrix theory. Our main technical contribution is to extend the resolvent technique to the fine-grained bias-variance characterization via a symmetrization argument (see Section B.3). Moreover, we emphasize that unlike [73, 62, 36], our result is non-asymptotic and does not require convergence of empirical spectral distributions.
3.3 Double Divergence under Ground-truth Featurization
@@ 4 Downstream-Optimal Feature Transfer
In this section, we explore the effects of upstream pretraining on downstream performance (Definition 2.2). Specifically, we optimize the average risk for an ensemble of downstream tasks with respect to the pretrained representation. We assume throughout that the ground-truth features $\mathbf{B}^{\star}$ and data covariance $\mathbf{\Sigma}$ are known. The motivation for this assumption is to detach the estimation effects of $\mathbf{B}^{\star}$ and $\mathbf{\Sigma}$ from their role in determining the optimal choice of $\widehat{\mathbf{B}}$. That said, we still include in Appendix H a sample-complexity analysis under the setting of Example 2.1 where $\mathbf{B}^{\star}$ is not known. Recall that in the sample-rich regime, $R$ is characterized by the limiting quantity $\sigma^{2}\frac{h}{n-h}$, regardless of $\widehat{\mathbf{B}}$. We thus focus on the sample-deficient regime where $n<h$ in this section.
4.1 An End-to-end Approach for Downstream Optimality
Below, we compare three approaches: (i) ignoring the pretrained features altogether and simply using the standard predictor, which can be obtained by $\lambda\to 0$; (ii) setting $\widehat{\mathbf{B}}=\mathbf{B}^{\star}$ and optimizing $(\lambda,\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}})$; and (iii) directly optimizing the pretraining objective to choose $\widehat{\mathbf{B}}$ and $(\lambda,\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}})$ using the results of Section 3. We then interpret the solution of approach (iii) to reveal the structure of optimal pretraining.
Definition 4.1 (Ridgeless predictor (RP).).
The method has no hyperparameters and sets
$\hat{y}(\mathbf{x}_{\mathrm{new}})={\mathbf{X}}^{\top}({\mathbf{X}}{\mathbf{X}}^{\top})^{+}\mathbf{y}.$
We include RP approach as a baseline, and it is expected to underperform when the available downstream data $n$ are limited. It is only optimal when $n\gg p$, where no dimension reduction is necessary
Definition 4.2 (Oracle-featurization predictor (OFP).).
The oracle-featurization predictor is the predictor $\hat{y}$ defined in Definition 2.2 with $\widehat{\mathbf{B}}=\mathbf{B}^{\star}$, i.e. the ground-truth featurization, and with $(\lambda,\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}})$ set optimally.
When $\mathbf{B}^{\star}$ is low-rank and the downstream data are limited (i.e. $q\ll n<p$), the ground-truth featurization can markedly increase performance by leveraging the shared representation. Counterintuitively, the OFP is not necessarily optimal. An example is when ${\bm{\alpha}^{\star}}$ is higher dimensional compared to ${{\bm{\beta}^{\star}}}$, i.e. $q>p$. Then estimating ${\bm{\alpha}^{\star}}$ first using $\hat{\bm{\alpha}}$ in Eq. (3) may incur larger risk than directly estimating ${{\bm{\beta}^{\star}}}$. Clearly, this example is outside the scope normally considered by transfer learning, since there is no shared low-dimensional structure. A more important reason the OFP is suboptimal is that it ignores the distribution of the covariates. Ideally, we would exploit structure in both $\mathbf{B}^{\star}$ and $\mathbf{\Sigma}$ when choosing $\widehat{\mathbf{B}}$. But, this is typically impossible either because the downstream task are not known at the time of pretraining or because the upstream-downstream procedure is not end-to-end differentiable. However, if we further assume that the task-specific paramter $\bm{\alpha}^{\star}$ are distributed i.i.d. from a prior $P_{\bm{\alpha}^{\star}}$ with zero mean and covariance $q^{-1}\mathbf{\Sigma}_{\bm{\alpha}^{\star}}\in\mathbb{R}^{q\times q}$, our theoretical results from Section 3.2 give us an analytic expression for the asymptotic behavior of downstream risk averaged across potential downstream tasks $R^{\mathsf{avg}}:=\mathbb{E}_{\bm{\alpha}^{\star}}R$:
$\mathfrak{R}^{\mathsf{avg}}:=\mathbb{E}_{\bm{\beta}^{\star}}\mathfrak{R}=\mathfrak{B}^{\mathsf{avg}}+\mathfrak{B}^{\mathsf{avg}}\mathcal{V}+\sigma^{2}\mathcal{V},$ (15)
@@ 4.1 An End-to-end Approach for Downstream Optimality
$\mathfrak{B}^{\mathsf{avg}}:=\mathbb{E}\mathfrak{B}\quantity(\widehat{\mathbf{B}},{\bm{\alpha}^{\star}},\mathbf{B}^{\star})=\frac{1}{q}\sum_{i\in H}\frac{\hat{t}_{i}\cdot\hat{\mathbf{w}}_{i}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\mathbf{B}^{\star}\mathbf{\Sigma}_{\bm{\alpha}^{\star}}{\mathbf{B}^{\star}}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\hat{\mathbf{w}}_{i}}{\left(1+\hat{t}_{i}b_{0}\right)^{2}}.$
@@
Remark 4.4 (A Backpropagation Routine).
To optimize $\mathfrak{R}^{\mathsf{avg}}$, we develop a custom backpropagation routine in Appendix G where we also discuss the non-standard step of backpropagating through the fixed point $b_{0}$.
@@
A minimax procedure may be developed such that one seeks to control the risk for the downstream risk for the worst task $R^{\mathrm{worst}}:=\max_{\norm{\bm{\alpha}^{\star}}_{2}^{2}\leq\mathfrak{c}}R,\mathfrak{c}>0$. Our procedure remains the same except that $\mathfrak{R}^{\mathrm{avg}}$ in Definition 4.3 is replaced with asymptotic limit of $\mathfrak{R}^{\mathsf{worst}}$ of $R^{\mathrm{worst}}$
$\mathfrak{R}^{\mathsf{worst}}:=\max_{{\bm{\alpha}^{\star}}\in\mathbb{B}^{q}(\sqrt{\mathfrak{c}})}\mathfrak{R}=\sigma^{2}\mathcal{V}+\mathfrak{c}\cdot\left(\mathcal{V}+1\right)\cdot\sigma_{\mathrm{max}}\quantity(\sum_{i\in H}\frac{\hat{t}_{i}\cdot{\mathbf{B}^{\star}}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\hat{\mathbf{w}}_{i}\hat{\mathbf{w}}_{i}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\mathbf{B}^{\star}}{\left(1+\hat{t}_{i}b_{0}\right)^{2}}),$
Note that $\mathfrak{R}^{\mathsf{worst}}$ remains end-to-end differentiable. See Appendix D and G for derivation of the above formula, theoretical guarantee and optimization routines.
Remark 4.6 (Methodological Contributions and Novelty).
Prior and concurrent works [73, 62, 36] have explored similar ideas of optimizing the generalized ridge penalty against the downstream risk. We remark on our methodological contributions and novelty in comparison to these works: prior works [73, 62] require the strong assumption that the penalty matrix $\hat{\bm{\Gamma}}$ commutes with data covariance $\mathbf{\Sigma}$. The concurrent work [36] dispenses the commutativity assumption but their optimization result only applies to a particular choice of ridge penalty strength (see [36], Theorem 2.5). We note that every additive component in the objective (15) corresponds to an FG bias or variance component; one can reweigh their contributions to adjust the bias-variance trade-off. Finally, our approach naturally leads to the minimax variant discussed in Remark 4.5, which is not considered in prior works.
Proposition 4.7 below supports the use of the objective $\mathfrak{R}^{\mathsf{avg}}$, as it implies that with a slightly stronger assumption on the prior distribution of $\bm{\alpha}^{\star}$, the risk for specific downstream tasks concentrates to $\mathfrak{R}^{\mathsf{avg}}$. See Appendix C for a proof.
@@
for the large constant $M$ in 3.1, where $\xi$ is a random vector with independent, sub-Gaussian, zero-mean, unit-variance entries and the sub-Gaussian norm bounded by $M$. Then, for any $D>0$, there exists $C=C(M,D)$ such that with probability at least $1-C(n^{-D}+q^{-D})$,
@@
4.2 Ablation Studies
Surprisingly, we find that the EEP easily finds $\widehat{\mathbf{B}}$ with lower risk than the RP and OFP. From Figure 2, we see that the EEP does this by making a better tradeoff of bias and variance. This is despite either the bias or variance of the other predictors being lower in some settings. In (d), the EEP is unbiased when $q<n$ by exploiting $\mathbf{B}^{\star}$’s low-rank structure, just like the OFP. Adjusting $\bm{\lambda}$ in both the OFP and the EEP removes the divergence in variance (compare (a) and (d) to Figure 1). However, Theorem 3.3 predicts that variance diverges when $h=n$ regardless of $\widehat{\mathbf{B}}$. Even when $\mathbf{\Sigma}$ includes a jitter term so that it is actually full-rank and only approximately low-rank (see (b) and (e)), the OFP still sees an explosion in variance as $n$ approaches the approximate rank. In contrast, the EEP is free to align its eigenvalues and eigenvectors to those of $\mathbf{\Sigma}$ and effectively modulate the variance at the expense of a slightly larger bias.
Finally, we see in (c) and (f), how as the capacity of $\widehat{\mathbf{B}}$ increases, its performance starts by matching that of the RP and eventually matches then exceeds that of the OFP. Notably, when $n>q$ (see (e) and (f)), OFP and EEP can completely remove the bias, which is not true when $n<q$ (see (d)); this corresponds to a phase transition between the hard- and soft-selection regime discussed in the next section. Section I.2 further demonstrates this by reproducing Figure 2 with different choices of $q$. See Section I.1 for detailed comparison of fine-grained risk components (i.e. $B,V_{\mathbf{X}},V_{\mathbf{X},{\bm{\varepsilon}}},V_{{\bm{\varepsilon}}}$) of these predictors, and Section I.3 for plots ablating other model parameters ($q$, SNR, and common structure in $\mathbf{B}^{\star}$).

### v2

@@ 3 Analytic Results for Downstream Risk
Recall that we introduced a linear predictor $\hat{y}=\hat{y}_{(\widehat{\mathbf{B}},\mathbf{y},\mathbf{X})},$ which leverages a pretrained featurization $\widehat{\mathbf{B}}$ and is fitted using data $(\mathbf{y},\mathbf{X})$ from a downstream task. We now derive asymptotic expressions for bias-variance decomposition of the risk of $\hat{y}$, defined as
$R:=\mathbb{E}_{{y}_{\mathrm{new}},\mathbf{x}_{\mathrm{new}}}\left({y}_{\mathrm{new}}-\hat{y}(\mathbf{x}_{\mathrm{new}})\right)^{2}.$
We emphasize the importance of fine-grained bias-variance decomposition because explicitly separating these sub-components of risk provides a clearer understanding of how different design choices in feature selection influence overall risk. In particular, we may investigate which $\widehat{\mathbf{B}}$ optimizes bias and variance component respectively. In Section 5, we attribute the observed sparsity in optimal featurization (i.e., feature selection) to the fine-grained bias and precisely characterize the associated phase transition phenomenon.
In this section, we treat $\widehat{\mathbf{B}}\in\mathbb{R}^{p\times k}$ as a fixed, deterministic input as we derive asymptotic expressions for risk and its bias-variance decompositions.
The classical bias-variance decomposition of the risk $\mathbb{E}_{{\bm{\varepsilon}}}R$ is defined as $B_{\mathsf{SC}}+V_{\mathsf{SC}}$ with
@@
for training data $\mathbf{y}=\mathbf{X}{\bm{\beta}^{\star}}+{\bm{\varepsilon}}$ and new response ${y}_{\mathrm{new}}={\mathbf{x}_{\mathrm{new}}}^{\top}{{\bm{\beta}^{\star}}}$. The bias-variance decomposition above is widely recognized in the statistics literature [24, 32, 23, 69]. However, as pointed out in [4], the above is conditional on $\mathbf{X}$ and only decomposes randomness in label noise ${\bm{\varepsilon}}$. Following [4], we consider the fine-grained bias-variance decomposition
@@
where the bias component is $B:=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\left({y}_{\mathrm{new}}-\mathbb{E}_{\mathbf{X}\!\!,{\bm{\varepsilon}}}\hat{y}\right)^{2}$ and variance component is $V:=V_{\mathbf{X}}+V_{\mathbf{X},{\bm{\varepsilon}}}+V_{{\bm{\varepsilon}}}$ with a further decomposition
@@
Here, the bias-variance decomposition is with respect to randomness in both the data and label noise; $V_{\mathbf{X}}$, $V_{{\bm{\varepsilon}}}$, and $V_{\mathbf{X},{\bm{\varepsilon}}}$ are a two-way ANOVA decomposition of $V$: $V_{\mathbf{X}}$ and $V_{{\bm{\varepsilon}}}$ are the variances explained by data and noise individually, and $V_{\mathbf{X},{\bm{\varepsilon}}}$ the additional variance explained by data and noise jointly. Unlike the classical decomposition, the fine-grained decomposition provides a clear interpretation of each additive component within the asymptotic risk formula (see (11) and discussion below). Additionally, as observed in [4] and [72] interpreting the classical bias can be challenging, especially as it tends to diverge at the interpolation boundary. The fine grained decomposition resolves this by attributing the divergence to the variance component $V_{\mathbf{X}}$, while the fine grained bias $B$ exhibits straightforward monotonic trends in our ablation studies (see Figure 1 and 2).
@@
Define $H:=\{i\in\{1,\ldots,p\}:\hat{t}_{i}\neq 0\}$, where |$\absolutevalue{H}=h=\mathrm{rank}(\mathbf{\Sigma})$. Let $b_{0}\in\mathbb{R}_{+}$ be the unique non-negative solution of
@@
The following theorem states that in the sample-deficient regime ($n<h=\mathrm{rank}(\mathbf{\Sigma})$), the total risk $R$ and its components $B,V_{\mathbf{X}},V_{\mathbf{X},{\bm{\varepsilon}}}$ converge to $\mathfrak{R}$ and $\mathfrak{B},\mathcal{V}\mathfrak{B}$ and $\mathcal{V}$ respectively whereas in the sample-rich regime ($n>h=\mathrm{rank}(\mathbf{\Sigma})$), the total risk has only one non-zero component $V_{\mathbf{X},{\bm{\varepsilon}}}$ and it converges to $\mathcal{U}$. We defer the proof to Appendix 6.1.
@@ 4 Downstream-Optimal Feature Transfer
In this section, we optimize the average risk across an ensemble of downstream tasks with respect to the pretrained feature matrix $\widehat{\mathbf{B}}$. In the sample-rich regime, the risk is determined by the limiting expression $\sigma^{2}\frac{h}{n-h}$, where $h=\operatorname{rank}(\mathbf{\Sigma})$, independent of $\widehat{\mathbf{B}}$. Therefore, our focus here is on the sample-deficient regime where $n<h$.
Setting the feature matrix to the ground truth, i.e., $\widehat{\mathbf{B}}=\mathbf{B}^{\star}$, is not necessarily optimal. Intuitively, the optimal $\widehat{\mathbf{B}}$ should account for both the distribution of covariates and the characteristics of the downstream task. A principled approach to selecting $\widehat{\mathbf{B}}$ would be to directly minimize the downstream risk. However, this is often impractical, either because downstream tasks are unknown at the time of pretraining or because the upstream-downstream process is not end-to-end differentiable. However, if we further assume that the task-specific parameter $\bm{\alpha}^{\star}$ are distributed i.i.d. from a prior $P_{\bm{\alpha}^{\star}}$ with zero mean and covariance $q^{-1}\mathbf{\Sigma}_{\bm{\alpha}^{\star}}\in\mathbb{R}^{q\times q}$, our theoretical results from Section 3 give us an analytic expression for the asymptotic behavior of downstream risk averaged across potential downstream tasks $R^{\mathsf{avg}}:=\mathbb{E}_{\bm{\alpha}^{\star}}R$:
$\mathfrak{R}^{\mathsf{avg}}:=\mathbb{E}_{\bm{\alpha}^{\star}}\mathfrak{R}=\mathfrak{B}^{\mathsf{avg}}+\mathfrak{B}^{\mathsf{avg}}\mathcal{V}+\sigma^{2}\mathcal{V},$ (12)
@@
$\mathfrak{B}^{\mathsf{avg}}:=\mathbb{E}_{\bm{\alpha}^{\star}}\mathfrak{B}\quantity(\widehat{\mathbf{B}},{\bm{\alpha}^{\star}},\mathbf{B}^{\star})=\frac{1}{q}\sum_{i\in H}\frac{\hat{t}_{i}\cdot\hat{\mathbf{w}}_{i}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\mathbf{B}^{\star}\mathbf{\Sigma}_{\bm{\alpha}^{\star}}{\mathbf{B}^{\star}}^{\top}\hat{\bm{\Gamma}}^{\frac{1}{2}}\hat{\mathbf{w}}_{i}}{\left(1+\hat{t}_{i}b_{0}\right)^{2}}.$
@@
A minimax procedure may be developed such that one seeks to control the risk for the downstream risk for the worst task $R^{\mathrm{worst}}:=\max_{\norm{\bm{\alpha}^{\star}}_{2}^{2}\leq\mathfrak{c}}R,\mathfrak{c}>0$. See Appendix C.
Proposition 4.3 below supports the use of the objective $\mathfrak{R}^{\mathsf{avg}}$, as it implies that with a slightly stronger assumption on the prior distribution of $\bm{\alpha}^{\star}$, the risk for specific downstream tasks concentrates to $\mathfrak{R}^{\mathsf{avg}}$. See Appendix B for a proof.
@@
for the constant $M$ in 3.1, where $\xi$ is a random vector with independent, sub-Gaussian, zero-mean, unit-variance entries and the sub-Gaussian norm bounded by $M$. Then, for any $D>0$, there exists $C=C(M,D)$ such that with probability at least $1-C(n^{-D}+q^{-D})$,
@@
Below and in Figure 2, we compare the end-to-end predictor (EEP) above with two other approaches with a series of ablation studies: (i) Ridgeless Predictor (RP): ignoring the pretrained features altogether and simply using the standard predictor, $\hat{y}(\mathbf{x}_{\mathrm{new}})={\mathbf{X}}^{\top}({\mathbf{X}}{\mathbf{X}}^{\top})^{+}\mathbf{y}$; (ii) Oracle-Featurization Predictor (OFP): setting $\widehat{\mathbf{B}}=\mathbf{B}^{\star}$ and optimizing $(\lambda,\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}})$.
We find that the EEP easily finds $\widehat{\mathbf{B}}$ with lower risk than the RP and OFP. From Figure 2, we see that the EEP does this by making a better tradeoff of bias and variance. This is despite either the bias or variance of the other predictors being lower in some settings. In (d), the EEP is unbiased when $q<n$ by exploiting $\mathbf{B}^{\star}$’s low-rank structure, just like the OFP. Adjusting $\bm{\lambda}$ in both the OFP and the EEP removes the divergence in variance (compare (a) and (d) to Figure 1). However, Theorem 3.3 predicts that variance diverges when $h=n$ regardless of $\widehat{\mathbf{B}}$. Even when $\mathbf{\Sigma}$ includes a jitter term so that it is actually full-rank and only approximately low-rank (see (b) and (e)), the OFP still sees an explosion in variance as $n$ approaches the approximate rank. In contrast, the EEP is free to align its eigenvalues and eigenvectors to those of $\mathbf{\Sigma}$ and effectively modulate the variance at the expense of a slightly larger bias.
Finally, we see in (c) and (f), how as the capacity of $\widehat{\mathbf{B}}$ increases, its performance starts by matching that of the RP and eventually matches then exceeds that of the OFP. Notably, when $n>q$ (see (e) and (f)), OFP and EEP can completely remove the bias, which is not true when $n<q$ (see (d)); this corresponds to a phase transition between the hard- and soft-selection regime discussed in the next section. Section G.2 further demonstrates this by reproducing Figure 2 with different choices of $q$. See Section G.1 for detailed comparison of fine-grained risk components (i.e. $B,V_{\mathbf{X}},V_{\mathbf{X},{\bm{\varepsilon}}},V_{{\bm{\varepsilon}}}$) of these predictors, and Section G.3 for a series of ablation studies on other model parameters ($q$, SNR, and common structure in $\mathbf{B}^{\star}$).